OpenStudy (compassionate):

8. During a 60 minute television show, the entertainment portion lasted 15 minutes less than 4 times the advertising portion. How long was the entertainment portion? @Miracrown

OpenStudy (compassionate):

60 = 15 - 4x 45 = 4x x = 11.5

OpenStudy (compassionate):

No wait: 60 = 4x - 15 75 = 4x x = 18.75

OpenStudy (nincompoop):

correct set up

OpenStudy (nincompoop):

it's 4x - 15 good that you caught it

OpenStudy (compassionate):

Was I correct?

OpenStudy (nincompoop):

100% correct

OpenStudy (compassionate):

i have a few questions about fractions and decimals. Up to answer them?

OpenStudy (nincompoop):

guide only, you know this laughing out loud

OpenStudy (nincompoop):

post them and I will determine how to best teach you

OpenStudy (compassionate):

wow u listen here

OpenStudy (compassionate):

im on a mobile device so forgiv speling g okay so if i ave 0.009 i can make it into a fraction by moving the decimal three places and putting it over o100 so 9/100 yes?

OpenStudy (nincompoop):

oh laughing out loud no two zeroes in the denominator tells you how many decimal points

Miracrown (miracrown):

9/1000

OpenStudy (nincompoop):

0.006 would be 6/1000

OpenStudy (compassionate):

is it becus the 9 is in the thousands place

Miracrown (miracrown):

@nincompoop #rude

OpenStudy (nincompoop):

the relationship of decimal and fraction

Miracrown (miracrown):